OpenStudy (anonymous):

Which of the following is not a step in solving 5x – 2(3x + 5) = 32?

OpenStudy (anonymous):

options?

OpenStudy (anonymous):

A. –x = 42 B. –1x = 42 C. –x + 10 = 32 D. 5x – 6x – 10 = 32

OpenStudy (anonymous):

C. 5x-2(3x+5)=32 5x-6x-10=32 -x-10=32 -x=42 or -1x=42 x=-42

OpenStudy (anonymous):

thanks for the medal @brianayala

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

do you understand it or need a better explanation?

OpenStudy (anonymous):

@KlOwNlOvE your welcome..

OpenStudy (anonymous):

i Understand, Thanks...

OpenStudy (anonymous):

you're welcome